OpenStudy (anonymous):

Find dy for y= sqrt(x^2 +4)

OpenStudy (anonymous):

you mean find dy/dx (derivative) ??

OpenStudy (anonymous):

No it says dy ... idk if its a misprint. That's why I'm really conused.

OpenStudy (anonymous):

they probably mean dy/dx (the first derivative). Do you know how to solve for that?

hartnn (hartnn):

find dy/dx whatever you get, multiply it by dx to get dy

OpenStudy (anonymous):

@hartnn is this implicit differentiation by any chance?

hartnn (hartnn):

you can do this without knowing implicit diff

OpenStudy (anonymous):

Yes and second derivative stuff

hartnn (hartnn):

this mainly requires chain rule, you know what is it ?

OpenStudy (anonymous):

Yes I do

hartnn (hartnn):

so, what is derivative of sqrt x ?

OpenStudy (anonymous):

@hartnn whats the distinguishing difference of 'dy' apposed to 'dx'? and not dy/dx?

hartnn (hartnn):

difference ? dy is a small change in 'y' dx is a small change in 'x' dy/dx is the slope of curve y basically you can separate the differentials. dy/dx =....., then dy = ..... dx

hartnn (hartnn):

jazzie, if you know derivative of sqrt x and chain rule, you'll be alble to solve this.

hartnn (hartnn):

*able

OpenStudy (anonymous):

Hey, @hartnn ! I'm attempting to solve this problem now. What would my steps be? y=(x^2+4)^(1/2) dy/dx = (1/2)(2x)(x^2+4)^(-1/2) dy/dx = x/[(x^2+4)^1/2] dy = x/[(x^2+4)^1/2] dx Is that it?

hartnn (hartnn):

yup! good work :)

OpenStudy (anonymous):

Thanks! :D